# Differential Mamba

### v1

@@ Abstract
Sequence models like Transformers and RNNs often overallocate attention to irrelevant context, leading to noisy intermediate representations. This degrades LLM capabilities by promoting hallucinations, weakening long-range and retrieval abilities, and reducing robustness. Recent work has shown that differential design can mitigate this issue in Transformers, improving their effectiveness across various applications. In this paper, we explore whether these techniques, originally developed for Transformers, can be applied to Mamba, a recent architecture based on selective state-space layers that achieves Transformer-level performance with greater efficiency. We show that a naive adaptation of differential design to Mamba is insufficient and requires careful architectural modifications. To address this, we introduce a novel differential mechanism for Mamba, empirically validated on language modeling benchmarks, demonstrating improved retrieval capabilities and superior performance over vanilla Mamba. Finally, we conduct extensive ablation studies and empirical analyses to justify our design choices and provide evidence that our approach effectively mitigates the overallocation problem in Mamba-based models. Our code is publicly available.
@@ 1 Introduction
Designing enhanced architectures for deep sequence modeling is a pivotal task in the ML community, as sequence models drive breakthroughs such as ChatGPT [1] in NLP and Stable diffusion [32] in computer vision, which are fundamental to modern generative models. However, these models face two major challenges as: efficiency, constrained by the quadratic time complexity in sequence length of the Transformer architecture [36], and robustness, which is hindered by inconsistency, reliability issues, and hallucinations, leading to sub-optimal performance. Our work addresses these challenges by enhancing the robustness of recent efficient architectures with sub-quadratic complexity such as Mamba [16], making them more reliable and robust.
To address this robustness issue, we aim to reduce the over-allocation of attention to irrelevant context across hidden layers, which often leads to noisy representations. Our approach is inspired by Ye et al. [43], who mitigated this problem in transformers through differential design. This architectural modification was the core of the Diff-Transformer, a transformer variant that demonstrated improved empirical performance, including greater robustness, enhanced retrieval and long-context capabilities compared to the original model.
Our focus is on improving the recently introduced Mamba architecture, which builds upon selective state-space layers (S6). This architecture is known for its efficiency, introducing sub-quadratic complexity in sequence length while also enabling auto-regressive decoding with a complexity that remains independent of sequence length. Beyond these efficiency advantages, recent studies have demonstrated that Mamba-based architectures can match or even surpass the SoTA performance of transformers, even at relatively large scales. For example, Falcon-Mamba 7B [46], a pure Mamba model, matches the performance of LLaMA- 3-8B on several language tasks. Additional notable examples of Mamba’s integration in LLMs include Jamba [23], Zamba [14], Hymba [11], and others [39, 37, 31]. More importantly, recent Mamba-based models have demonstrated remarkable performance as reasoning models [28, 40], underscoring their central role in the ongoing test-time scaling revolution, as exemplified by models such as OpenAI’s O1 [20] and DeepSeek’s R1 [19].
@@
where we hypothesize that improving robustness by mitigating the over-allocation problem can enhance the model’s general capabilities, improve retrieval and long-context processing, and potentially reduce issues related to consistency and hallucinations in LLMs.
@@
Our main contributions are as follows: (i) We present Diff-Mamba, a modification of the Mamba architecture inspired by Diff-Transformer, which mitigates the problem of over-allocating attention scores to irrelevant context and improves the general language modeling abilities of the model. (ii) Through a series of ablation studies and empirical analysis using mechanistic interpretability tools, we justify our design choices and demonstrate that the intermediate representations obtained from our method are indeed less noisy. (iii) Finally, we show that Diff-Mamba demonstrates improved retrieval and long-context capabilities compared to Mamba. This is particularly important, as recurrent LLMs such as Mamba are primarily designed to address the inefficiencies caused by the quadratic complexity of transformers, which becomes especially critical in long-context regimes. Accordingly, Diff-Mamba achieves improvements over Mamba precisely in the domain where such architectures are most needed [6].
@@ 2 Background
Here we describe the scientific context and introduce the key terminology for discussing our method.
@@ 2.1 Differential Transformer
Self-Attention is a fundamental component of Transformer architectures [36], has significantly shaped recent advances in both NLP and computer vision. This mechanism enables dynamic focus allocation by capturing pairwise token dependencies, allowing the model to determine the relative importance of each token within a sequence. Mathematically, it is expressed as:
$Attention(Q,K,V)=\alpha V,\quad\alpha=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)$ (1)
@@
To address the problem of over-allocation of attention to irrelevant tokens, Ye et al. [43], introduced Diff Transformer, a mechanism that reduces attention noise through differential denoising by splitting each attention head into two and subtracting one attention map from the other. This mechanism can be described as follows:
$DiffAttention(Q_{1},K_{1},Q_{2},K_{2},V)=(\alpha_{1}-\lambda\alpha_{2}),\quad\alpha_{i}=\text{softmax}\left(\frac{Q_{i}{K_{i}}^{T}}{\sqrt{d_{k}}}\right)$ (2)
where $\lambda$ is is a learnable scalar. To better improve the training dynamics, $\lambda$ is re-parameterized and Group Normalization [42] is applied at the end of each head (post-subtraction).
@@ 2.2 State-Space Layers
State-space layers were first introduced in Gu et al. [17] and were later substantially improved by the S4 model [18]. Since then, they have demonstrated strong performance across a variety of domains, including NLP [12, 26], audio generation [15], image modeling [3, 27], long-horizon video understanding [38], reinforcement learning [9, 25], and speech recognition [33]. These models implement linear recurrent update rules derived from time-invariant state-space formulations, which can be efficiently computed in parallel using convolutions and with sub-quadratic complexity.
@@ 2.3 Mamba and Selective State-Space Layers
$X=\sigma(\text{Conv1D}(\text{Linear}(U)),\quad Z=\sigma(\text{Linear}(U)),\quad Y=\text{S6}(X),\quad\hat{Y}=\text{Linear}(Y\otimes Z)$ (3)
@@
$h_{t}=\bar{A}h_{t-1}+\bar{B}x_{t},\quad y_{k}=Ch_{t},\quad\bar{A}=f_{A}(A,\Delta),\quad\bar{B}=f_{B}(B,\Delta)$ (4)
where $f_{A},f_{B}$ are discretization functions, and the discrete system matrices are $\bar{A}\in\mathbb{R}^{N^{\prime}\times N}$ and $\bar{B}\in\mathbb{R}^{N^{\prime}\times 1}$. The recurrent rule in Eq. 4 can be computed efficiently in parallel on modern hardware accelerators using work-efficient parallel scans [35] or a simple scalar convolution via FFTs [17]. Note that Eq. 4 is a map from $\mathbb{R}^{L}$ to $\mathbb{R}^{L}$, and to process $D$ channels, multiple independent instances are used.
The S6 layer differs from standard SSMs by employing a selective mechanism, where the system matrices are input-dependent. As a result, the system becomes time-invariant, with the per-step system matrices determined by the entire set of channels and then applied to process each channel independently. The entire mechanism can be computed by: $S_{B},S_{C}\in\mathbb{R}^{N^{\prime}\times D}$, $A\in\mathbb{R}^{D\times N^{\prime}}$ and $S_{\Delta}\in\mathbb{R}^{1\times D}$ to define the time-variant matrices as follows:
$B_{t}=S_{B}X_{*t},\ C_{t}=S_{C}X_{*t},\ \Delta_{t}=\text{softplus}(S_{\Delta}X_{*t}),\quad\bar{A}_{t}=\exp(\Delta_{t}A),\quad\bar{B}_{t}=\Delta_{t}B_{t}$ (5)
@@ 2.3.1 Mamba as Implicit Attention
The connection between Mamba and linear attention layers is well-established in [2, 8, 34]. Specifically, it has been shown that the time-variant recurrent update rule of S6 for a single channel (see Eq. 6) can be explicitly unrolled into the linear attention formulation $Y=AX$ when $A$ is an implicit attention matrix defined by:
$\begin{bmatrix}C_{1}\bar{B}_{1}&0&\cdots&0\\ C_{2}\bar{A}_{2}\bar{B}_{1}&C_{2}\bar{B}_{2}&\cdots&0\\ \vdots&\vdots&\ddots&0\\ C_{L}\Pi_{k=2}^{L}\bar{A}_{k}\bar{B}_{1}\quad&C_{L}\Pi_{k=3}^{L}\bar{A}_{k}\bar{B}_{2}\quad&\cdots\quad&C_{L}\bar{B}_{L}\end{bmatrix}$ (7)
This perspective is further extended by Zimerman et al. [45], who generalize the interpretation of S6 as implicit attention from Eq. 7 to encompass most components of the Mamba block, including activations, normalization layers, convolutional layers, and the gate branch, into a unified implicit attention formulation.
@@
The formulation of data-controlled linear operators was first introduced by Poli et al. [30], who demonstrated that self-attention can be viewed as an expressive form of such operators. This principle guided the authors in designing the Hyena layer. Subsequently, Ali et al. [2] showed that S6 layers could also be unified under an implicit variant of this formulation, which Zimerman et al. [45] further extended to additional architectures, including the entire Mamba block, RWKV [29], and Griffin [10]. This extended perspective inspired our approach, leading us to interpret differential design as a method to implicitly parameterize less noisy data-controlled linear operators. This insight motivated our decision to apply differential design at the Mamba level.
@@ 3 Method
We begin with the simplest implementation of incorporating the differential mechanism into the Mamba block. Our approach is inspired by the Differential Transformer [43], which applies subtraction at the attention-level rather than at the transformer level. This mechanism is built on top of self attention and it can be described as:
@@
Diff $\text{Diff Attn}(x)=\Big{(}\mathcal{S}(\frac{Q_{1}K_{1}^{T}}{\sqrt{d}})-\lambda\mathcal{S}(\frac{Q_{2}K_{2}^{T}}{\sqrt{d}})\Big{)}V$ (9)
where $\mathcal{S}$ is the softmax function and $\lambda$ is uniquely parameterized to ensure stability and improve training dynamics:
@@
To better adapt this technique to Mamba models, we reinterpret differential attention through the lens of a data-controlled linear operator [30], leading to the following formulation:
$\text{Diff Attn}(x)=AV,\quad A=A_{1}-\lambda A_{2},\quad A_{1}=\mathcal{S}(\frac{Q_{1}K_{1}^{T}}{\sqrt{d}}),\quad A_{2}=\mathcal{S}(\frac{Q_{2}K_{2}^{T}}{\sqrt{d}})$ (11)
@@ 3.1 Diff S6
[…]
where $A_{1}$ and $A_{2}$ are the implicit attention matrices of S6 controlled by the system matrices $\bar{A}_{ij}$,$\bar{B}_{ij}$ and $C_{ij}$ for any time-step $j\in[L]$ and model index $j\in{1,2}$ defined as follows:
$\begin{bmatrix}C_{i1}\bar{B}_{i1}&0&\cdots&0\\ C_{i2}\bar{A}_{i2}\bar{B}_{i1}&C_{i2}\bar{B}_{i2}&\cdots&0\\ \vdots&\vdots&\ddots&0\\ C_{iL}\Pi_{k=2}^{L}\bar{A}_{ik}\bar{B}_{i1}\quad&C_{iL}\Pi_{k=3}^{L}\bar{A}_{ik}\bar{B}_{i2}\quad&\cdots\quad&C_{iL}\bar{B}_{iL}\end{bmatrix}$ (14)
One crucial difference between Diff S6 and Diff Attention (see Eqs. 12 and  9) is that Diff Attention subtracts elements on the same scale, as softmax produces values in the range [0,1]. In contrast, S6 produces unnormalized and unbounded outputs. To address this discrepancy, we introduce an additional normalization step denoted by $\mathbb{N}$:
-Diff $\mathbb{N}\text{-Diff S6}(X)=\mathbb{N}({S6}_{1}(X)-\lambda{S6}_{2}(X))$ (15)
@@
where $\bar{\lambda}$ is a learnable parameter $\bar{\lambda}\in\mathbb{R}^{D}$ used to parameterize $\lambda$ as a positive and more stable weight.

### v2

@@ Abstract
Sequence models like Transformers and RNNs often overallocate attention to irrelevant context, leading to noisy intermediate representations. This degrades LLM capabilities by promoting hallucinations, weakening long-range and retrieval abilities, and reducing robustness. Recent work has shown that differential design can mitigate this issue in Transformers, improving their effectiveness across various applications. In this paper, we explore whether these techniques, originally developed for Transformers, can be applied to Mamba, a recent architecture based on selective state-space layers that achieves Transformer-level performance with greater efficiency. We show that a naive adaptation of differential design to Mamba is insufficient and requires careful architectural modifications. To address this, we introduce a novel differential mechanism for Mamba, empirically validated on language modeling benchmarks, demonstrating improved retrieval, long-context capabilities and superior performance over vanilla Mamba. Finally, we conduct extensive ablation studies and empirical analyses to justify our design choices and provide evidence that our approach effectively mitigates the overallocation problem in Mamba-based models. Our code is publicly available.
@@
Differential Mamba
Nadav Schneider1,2, Itamar Zimerman3,4, Eliya Nachmani1 1School of Electrical and Computer Engineering, Ben-Gurion University of the Negev 2IAEC 3Tel-Aviv University 4IBM Research nadavsch@post.bgu.ac.ilzimerman1@mail.tau.ac.ileliyanac@bgu.ac.il
@@ 1 Introduction
Designing enhanced architectures for deep sequence modeling is a pivotal task in the ML community, as sequence models drive breakthroughs such as ChatGPT (Achiam et al., 2023) in NLP and Stable diffusion (Rombach et al., 2022) in computer vision, which are fundamental to modern generative models. However, these models face two major challenges as: efficiency, constrained by the quadratic time complexity in sequence length of the Transformer architecture (Vaswani, 2017), and robustness, which is hindered by inconsistency, reliability issues, and hallucinations, leading to sub-optimal performance. Our work addresses these challenges by enhancing the robustness of recent efficient architectures with sub-quadratic complexity such as Mamba (Gu and Dao, 2024), making them more reliable and robust.
To address this robustness issue, we aim to reduce the over-allocation of attention to irrelevant context across hidden layers, which often leads to noisy representations. Our approach is inspired by Ye et al. (2024), who mitigated this problem in transformers through differential design. This architectural modification was the core of the Diff-Transformer, a transformer variant that demonstrated improved performance, including greater robustness, enhanced retrieval and long-context capabilities compared to the original model.
Our focus is on improving the recently introduced Mamba architecture, which builds upon selective state-space layers (S6). This architecture is known for its efficiency, introducing sub-quadratic complexity in sequence length while also enabling auto-regressive decoding with a complexity that remains independent of sequence length. Beyond these efficiency advantages, recent studies have demonstrated that Mamba-based architectures can match or even surpass the SoTA performance of transformers, even at relatively large scales. For example, Falcon-Mamba 7B (Zuo et al., 2024), a pure Mamba model, matches the performance of LLaMA- 3-8B on several language tasks. Additional notable examples of Mamba’s integration in LLMs include Jamba (Lieber et al., 2024), Zamba (Glorioso et al., 2024), Hymba (Dong et al., 2024), and others (Wang et al., 2024; Waleffe et al., 2024; Ren et al., 2024). More importantly, recent Mamba-based models have demonstrated remarkable performance as reasoning models (Paliotta et al., 2025; Wang et al., 2025), underscoring their central role in the ongoing test-time scaling revolution, as exemplified by models such as OpenAI’s O1 (Jaech et al., 2024) and DeepSeek’s R1 (Guo et al., 2025).
@@
where we hypothesize that improving robustness by mitigating the over-allocation problem can enhance the model’s general capabilities, improve retrieval and long-context processing, and potentially reduce issues related to consistency and hallucinations.
@@
Our main contributions are as follows: (i) We present Diff-Mamba, a modification of the Mamba architecture inspired by Diff-Transformer, which mitigates the problem of over-allocating attention scores to irrelevant context and improves the general language modeling abilities of the model. (ii) Through a series of ablation studies and empirical analysis using mechanistic interpretability tools, we justify our design choices and demonstrate that the intermediate representations obtained from our method are indeed less noisy. (iii) Finally, we show that Diff-Mamba demonstrates improved retrieval and long-context capabilities compared to Mamba. This is particularly important, as recurrent LLMs such as Mamba are primarily designed to address the inefficiencies caused by the quadratic complexity of transformers, which becomes especially critical in long-context regimes. Accordingly, Diff-Mamba achieves improvements over Mamba precisely in the domain where such architectures are most needed (Ben-Kish et al., 2025b).
@@ 2 Background
Here we describe the scientific context and introduce the terminology for discussing our method.
@@ 2.1 Differential Transformer
Self-Attention is a fundamental component of Transformer architectures (Vaswani, 2017), has significantly shaped recent advances in both NLP and computer vision. This mechanism enables dynamic focus allocation by capturing pairwise token dependencies, allowing the model to determine the relative importance of each token within a sequence. Mathematically, it defined by:
$\text{Attn}(Q,K,V)=\alpha V,\quad\alpha=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)$ (1)
@@
To address the problem of over-allocation of attention to irrelevant tokens, Ye et al. (2024), introduced Diff-Transformer, a mechanism that reduces attention noise through differential denoising by splitting each attention head into two, and subtracting one attention map from the other. This mechanism is defined by:
$\text{DiffAttn}(Q_{1},K_{1},Q_{2},K_{2},V)=(\alpha_{1}-\lambda\alpha_{2})V$ (2)
$\alpha_{i}=\text{softmax}\left(\frac{Q_{i}{K_{i}}^{T}}{\sqrt{d_{k}}}\right)$ (3)
where $\lambda$ is a learnable scalar. To better improve the training dynamics, $\lambda$ is re-parameterized and Group Normalization (Wu and He, 2018) is applied at the end of each head (post-subtraction).
@@ 2.2 State-Space Layers
State-space layers were first introduced in Gu et al. (2021) and were later substantially improved by the S4 model (Gu et al., 2022). Since then, they have demonstrated strong performance across a variety of domains, including NLP (Fu et al., 2022; Mehta et al., 2022), audio generation (Goel et al., 2022), image modeling (Baron et al., 2023; Nguyen et al., 2022), long-horizon video understanding (Wang et al., 2023), reinforcement learning (David et al., 2022; Lu et al., 2024), and speech recognition (Saon et al., 2023). These models implement linear recurrent update rules derived from time-invariant state-space formulations, which can be efficiently computed in parallel using convolutions and with sub-quadratic complexity.
@@ 2.3 Mamba and Selective State-Space Layers
$\begin{split}X&=\sigma(\text{Conv1D}(\text{Linear}(U))),\quad Z=\sigma(\text{Linear}(U))\\ Y&=\text{S6}(X),\quad\hat{Y}=\text{Linear}(Y\otimes Z)\end{split}$ (4)
@@
$\begin{split}h_{t}&=\bar{A}h_{t-1}+\bar{B}x_{t},\quad y_{k}=Ch_{t}\\ \bar{A}&=f_{A}(A,\Delta),\quad\bar{B}=f_{B}(B,\Delta)\end{split}$ (5)
where $f_{A},f_{B}$ are discretization functions, and the discrete system matrices are $\bar{A}\in\mathbb{R}^{N^{\prime}\times N}$ and $\bar{B}\in\mathbb{R}^{N^{\prime}\times 1}$. The recurrent rule in Eq. 5 can be computed efficiently in parallel on modern hardware accelerators using work-efficient parallel scans (Smith et al., 2022) or a simple scalar convolution via FFTs (Gu et al., 2021). Note that Eq. 5 is a map from $\mathbb{R}^{L}$ to $\mathbb{R}^{L}$, and to process $D$ channels, multiple independent instances are used.
The S6 layer differs from standard SSMs by employing a selective mechanism, where the system matrices are input-dependent. As a result, the system becomes time-invariant, with the per-step system matrices determined by the entire set of channels and then applied to process each channel independently. The entire mechanism can be computed by: $S_{B},S_{C}\in\mathbb{R}^{N^{\prime}\times D}$, $A\in\mathbb{R}^{D\times N^{\prime}}$ and $S_{\Delta}\in\mathbb{R}^{1\times D}$ to define the time-variant matrices by:
$\begin{split}B_{t}&=S_{B}X_{*t},\;C_{t}=S_{C}X_{*t},\;\Delta_{t}=\text{softplus}(S_{\Delta}X_{*t})\\ \bar{A}_{t}&=\exp(\Delta_{t}A),\;\bar{B}_{t}=\Delta_{t}B_{t}\end{split}$ (6)
@@ 2.3.1 Mamba as Implicit Attention
The connection between Mamba and linear attention layers is well-established in (Ali et al., 2024; Dao and Gu, 2024; Sieber et al., 2025). Specifically, it has been shown that the time-variant recurrent update rule of S6 for a single channel (see Eq. 7) can be explicitly unrolled into the linear attention formulation $Y=AX$ when $A$ is an implicit attention matrix defined by:
$\begin{bmatrix}C_{1}\bar{B}_{1}&0&\cdots&0\\ C_{2}\bar{A}_{2}\bar{B}_{1}&C_{2}\bar{B}_{2}&\cdots&0\\ \vdots&\vdots&\ddots&0\\ C_{L}\!\!\prod_{k=2}^{L}\bar{A}_{k}\bar{B}_{1}&C_{L}\!\!\prod_{k=3}^{L}\bar{A}_{k}\bar{B}_{2}&\cdots&C_{L}\bar{B}_{L}\end{bmatrix}$ (8)
This perspective is further extended by Zimerman et al. (2024), who generalize the interpretation of S6 as implicit attention from Eq. 8 to encompass most components of the Mamba block, including activations, normalization layers, convolutional layers, and the gate branch, into a unified implicit attention formulation.
@@
The formulation of data-controlled linear operators was first introduced by Poli et al. (2023), who demonstrated that self-attention can be viewed as an expressive form of such operators. This principle guided the authors in designing the Hyena layer. Subsequently, Ali et al. (2024) showed that S6 layers could also be unified under an implicit variant of this formulation, which Zimerman et al. (2024) further extended to additional architectures, including the entire Mamba block, RWKV (Peng et al., 2023), and Griffin (De et al., 2024). This extended perspective inspired our approach, leading us to interpret differential design as a method to implicitly parameterize less noisy data-controlled linear operators. This insight motivated our decision to apply differential design at the Mamba level.
@@ 3 Method
We begin with the simplest implementation of incorporating the differential mechanism into the Mamba block. Our approach is inspired by the Differential Transformer (Ye et al., 2024), which applies subtraction at the attention-level rather than at the transformer level. This mechanism is built on top of self attention and it can be described as:
@@
$\text{DiffAttn}(x)=\Big(\mathcal{S}(\frac{Q_{1}K_{1}^{T}}{\sqrt{d}})-\lambda\mathcal{S}(\frac{Q_{2}K_{2}^{T}}{\sqrt{d}})\Big)V\,$ (10)
Here, $\mathcal{S}$ is the softmax and $\lambda$ is parameterized to ensure stability and improve training dynamics:
@@
To better adapt this technique to Mamba models, we reinterpret differential attention through the lens of a data-controlled linear operator (Poli et al., 2023), leading to the following formulation:
$\begin{split}\text{DiffAttn}(x)&=AV,\quad A=A_{1}-\lambda A_{2}\\ A_{1}&=\mathcal{S}\left(\frac{Q_{1}K_{1}^{T}}{\sqrt{d}}\right),\quad A_{2}=\mathcal{S}\left(\frac{Q_{2}K_{2}^{T}}{\sqrt{d}}\right)\end{split}$ (12)
@@ 3.1 Diff S6
$\text{Diff S6}(x)=(A_{1}-\lambda A_{2})X=AX$ (14)
where $A_{1}$ and $A_{2}$ are the implicit attention matrices of S6 controlled by the system matrices $\bar{A}_{ij}$$\bar{B}_{ij}$ and $C_{ij}$ for any time-step $j\in[L]$ and model index $j\in{1,2}$ defined as follows:
$\begin{bmatrix}C_{i1}\bar{B}_{i1}&0&\cdots&0\\ C_{i2}\bar{A}_{i2}\bar{B}_{i1}&C_{i2}\bar{B}_{i2}&\cdots&0\\ \vdots&\vdots&\ddots&0\\ C_{iL}\!\!\prod_{k=2}^{L}\bar{A}_{ik}\bar{B}_{i1}&C_{iL}\!\!\prod_{k=3}^{L}\bar{A}_{ik}\bar{B}_{i2}&\cdots&C_{iL}\bar{B}_{iL}\end{bmatrix}$ (15)
One crucial difference between Diff S6 and Diff Attention (see Eqs. 13 and  10) is that Diff Attention subtracts elements on the same scale, as softmax produces values in the range [0,1]. In contrast, S6 produces unnormalized and unbounded outputs. To address this discrepancy, we introduce an additional normalization step denoted by $\mathbb{N}$:
-$\mathbb{N}\text{-Diff S6}(X)=\mathbb{N}({S6}_{1}(X)-\lambda{S6}_{2}(X))$ (16)
@@
where $\bar{\lambda}\in\mathbb{R}^{D}$ is a learnable parameter used to parameterize $\lambda$ as a positive and more stable weight.